### v2

@@ 5.1 Simulation results
  • trade-off rule $\widehat{g}_{\text{tradeoff}}$ with trade-off coefficient $\bar{\lambda}$ described in Section 6.2.
@@
$\widehat{g}_{\text{sample}}$ $\widehat{g}_{\alpha}$ $\widehat{g}_{\text{tradeoff}}$
Prob. of selecting infeasible policies 35.4% 8% 79.6%
Prob. of selecting suboptimal policies 87.0% 98.6% 37.6%
Average welfare loss 0.06 0.60 -0.02
Average cost -\$3 -\$57 \$105
Notes: This table reports properties of statistical rules $\widehat{g}$, as averaged over 500 simulations. Row 1 reports the probability that the rule selects an eligibility policy that violates the budget constraint, i.e. $Pr_{P^{n}}\{B(\widehat{g};P)>0\}$. Row 2 reports the probability that the rule achieves strictly less welfare than the constrained optimal policy $g_{P}^{\ast}$, i.e. $Pr_{P^{n}}\{W(\widehat{g};P)<W(g_{P}^{\ast};P)\}$. Row 3 reports the average welfare loss of the rule relative to the maximum feasible welfare, i.e. $\frac{E_{P^{n}}\left[W(g_{P}^{\ast};P)-W(\widehat{g};P)\right]}{W(g_{P}^{\ast};P)}$. Row 4 reports the average cost of the policies selected by the rule, i.e. $E_{P^{n}}\left[B(\widehat{g};P)\right]$.
Row 1 of Table 5.1 illustrates that it is possible for all three statistical rule $\widehat{g}$ to select infeasible policies. A lower probability of selecting infeasible policies suggests the rule is closer to achieving asymptotic feasibility. In the distribution calibrated to the OHIE sample, the original sample-analog rule $\widehat{g}_{\text{sample}}$ might not be asymptotically feasible as it can select infeasible eligibility policies in 35.4% of the draws. In contrast, Theorem 3.1 guarantees that a simple modification $\widehat{g}_{\alpha}$ selects infeasible eligibility policies in less than 5% of the draws, regardless of the distribution. Simulation confirms such guarantee as the mistakes only happen 8% of the time.
@@
In the distribution calibrated to the OHIE, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ on average achieves higher welfare than the sample-analog rule $\widehat{g}_{\text{sample}}$. As shown in row 3 of Table 5.1, the welfare loss of $\widehat{g}_{\text{tradeoff}}$ is -2% of the maximum feasible welfare $W(g_{P}^{\ast};P)$, compared to 6% for $\widehat{g}_{\text{sample}}$. However, its improvement can be at the cost of violating the budget constraint more often than $\widehat{g}_{\text{sample}}$, at a rate of 79.6%. Though as shown in row 4 of Table 5.1, on average the violation is limited, which confirms Theorem 4.1.
6 Medicaid expansion: empirical illustration
Example 2.1 of Section 2 explains a more flexible Medicaid expansion policy that would allow the income thresholds to vary with the number of children. This section illustrates this example using data from the the Oregon Medicaid Health Insurance Experiment (OHIE).
6.1 Data
The original OHIE sample consists of 74,922 individuals (representing 66,385 households). Of these, 26,423 individuals responded to the initial mail survey, which collects information on income as percentage of the federal poverty level and number of children, which are the characteristics $X_{i}$ of interest for targeting. [^2] More accurately, I follow Sacarny et al. (2020) to approximate number of children by the number of family members under age 19 living in house as reported on the initial mail survey. I exclude individuals who did not respond to the initial survey from my sample, which differs from the sample analyzed in Finkelstein et al. (2012) as I focus on individuals who responded both to the initial and the main surveys from the OHIE. Due to this difference, the expansion policies selected using my sample do not directly carry their properties to the population underlying the original OHIE sample, as the distributions of $X$ differ. After one year, the main survey collects data related to health ($Y_{i}$), health care utilization ($C_{i}$) and actual enrollment in Medicaid ($M_{i}$), which allows me to construct estimates for the benefit and cost of Medicaid eligibility $(\Gamma,R)$. Therefore I further exclude individuals who did not respond to the main survey from my sample.
For health ($Y_{i}$), I follow the binary measurement in Finkelstein et al. (2012) based on self-reported health, where an answer of “poor/fair” is coded as $Y_{i}=0$ and “excellent/very good/good” is coded as $Y_{i}=1$. For health care utilization ($C_{i}$), the study collected measures of utilization of prescription drugs, outpatient visits, ER visits, and inpatient hospital visits. Finkelstein et al. (2012) annualize these utilization measures to turn these into spending estimates, weighting each type by its average cost (expenditures) among low-income publicly insured non-elderly adults in the Medical Expenditure Survey (MEPS). Note that health and health care utilization are not measured at the same scale, which requires rescaling when I consider the trade-off between the two. I address this issue in Section 6.2. Lastly, since the enrollment in Medicaid still requires an application, not everyone eligible in the OHIE eventually enrolled in Medicaid, which implies $M_{i}\leq D_{i}$.
Given the setup of the OHIE, Medicaid eligibility ($D_{i}$) is random conditional on household size (number of adults in the household) entered on the lottery sign-up form and survey wave. While the original experimental setup would ensure randomization given household size, the OHIE had to adjust randomization for later waves of survey respondents (see the Appendix of Finkelstein et al. (2012) for more details). Denote the confounders (household size and survey wave) with $V_{i}$, and define the propensity score as $p(V_{i})=Pr\{D_{i}=1\mid V_{i}\}$. If the propensity score is known, then the construction of the estimates follows directly from the formula (10). However, the adjustment for later survey waves means I need to estimate the propensity score, and I adapt the formula (10) following Athey and Wager (2021) to account for the estimated propensity score.
Specifically, define the conditional expectation function (CEF) of a random variable $U_{i}$ as $\gamma^{U}=E[U_{i}\mid V_{i},D_{i}$]. Since $V_{i}$ in my case is discrete, I use a fully saturated model to estimate the propensity score $\widehat{p}(V_{i})$ and the CEF $\widehat{\gamma}^{U}(V_{i},D_{i})$. I then form the estimated Horvitz-Thompson weight with the estimated propensity score as $\widehat{\alpha}(V_{i},D_{i})=\frac{D_{i}}{\widehat{p}(V_{i})}-\frac{1-D_{i}}{1-\widehat{p}(V_{i})}$. For health benefit due to Medicaid eligibility, define the estimate $\Gamma_{i}^{\ast}=\widehat{\gamma}^{Y}(V_{i},1)-\widehat{\gamma}^{Y}(V_{i},0)+\widehat{\alpha}(V_{i},D_{i})\cdot\left(Y_{i}-\widehat{\gamma}^{Y}(V_{i},D_{i})\right)$. The 2014 Medicaid spending was roughly \$6,000 per adult enrollee in Oregon, according to the expenditure information obtained from MACPAC (2019). To formalize the budget constraint that the per enrollee cost of the proposed policy cannot exceed the 2014 policy, I need to account for imperfect take-up because not everyone eligible for Medicaid would enroll. For the per enrollee excess cost relative to the current level, define the estimate $R_{i}^{\ast}=\widehat{\gamma}^{Z}(V_{i},1)+\frac{D_{i}}{\widehat{p}(W_{i})}\cdot\left(Z_{i}-\widehat{\gamma}^{Z}(V_{i},D_{i})\right)$ where $Z_{i}=C_{i}-\$6,000\cdot M_{i}$.
Table 6.1: Summary statistics of the OHIE sample by number of children
Number of children Sample size Sample mean of $\Gamma_{i}^{\ast}$ Sample mean of $R_{i}^{\ast}$
0 5,758 3.1% \$651
(0.01) (138)
1 1,736 10.3% \$348
(0.02) (247)
$\geq 2$ 2,641 1.5% -\$275
(0.02) (150)
Full sample 10,135 4.0% \$358
(0.01) (97)
Notes: This table presents summary statistics on the sample of individuals who responded to both the initial and the main surveys from the Oregon Health Insurance Experiment (the OHIE sample). The first three rows represent individuals living with different number of children (family members under age 19), and the last row is the aggregate. The estimate for benefit $\Gamma_{i}^{\ast}$ is an estimate for the increase in the probability of an individual reporting “excellent/very good/good” on self-reported health (as opposed to “poor/fair”) after receiving Medicaid eligibility. The estimate for cost $R_{i}^{\ast}$ is an estimate for individual’s health care expenditure that needs to be reimbursed by Medicaid, in excess of the current policy. Standard errors are shown in parentheses below.
Table 6.1 presents the summary statistics. While Online Appendix B.3 argues the estimation errors in $\Gamma_{i}^{\ast}$ and $R_{i}^{\ast}$ are asymptotically negligible, in finite samples, the cost estimates are highly variable, resulting in noisy estimate $\widehat{B}_{n}(g)$.
6.2 Budget-constrained Medicaid expansion
Figure 6.1 summarizes the selected expansion policies, which are income thresholds specific to the number of children. The sample-analog rule $\widehat{g}_{\text{sample}}$ chooses to restrict Medicaid eligibility, especially lowering the income threshold for childless individuals far below the current level. As a result, the budget estimate for the selected policy is -\$0.6, just below the threshold of zero. However, due to the large variation in the cost estimates as illustrated in Table 6.1 , meeting the sample budget constraint still involves uncertainty about whether the selected policy is welfare-efficient in the population as argued in Proposition 3.1.
To construct the trade-off rule $\widehat{g}_{\text{tradeoff}}$ as proposed in Section 4, I need to choose $\overline{\lambda}$, the upper bound on the marginal gain from violating the constraint. In my empirical illustration, the budget constraint is in terms of monetary value. The objective function, however, is measured based on self-reported health, which does not directly translate to a monetary value. Following Finkelstein et al. (2019), I convert self-reported health into value of a statistical life year (VSLY) based on existing estimates. Specifically, a conservative measure for the increase in quality-adjusted life year (QALY) when self-reported health increases from “poor/fair” to “excellent/very good/good” is roughly 0.6. The “consensus” estimate for the VSLY for one unit of QALY from Cutler (2004) is \$100,000 for the general US population. Taken these estimates together, I set $\overline{\lambda}=1/(0.6\cdot 10^{5})$. Given this upper bound, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ chooses to assign Medicaid eligibility to more individuals, and to raise the income thresholds above the current level for those with children. Therefore the budget estimate for the selected policy is \$110, slightly above the budget constraint. The higher level occurs because on average the benefit estimates are positive, and the trade-off rule finds that the additional health benefit from violating the budget constraint exceeds the cost of doing so, especially those with children. While a traditional cost-benefit analysis would also recommend prioritizing those with children based on Table 6.1, the trade-off rule offers a more interpretable recommendation by directly selecting income thresholds.
Figure 6.1: More flexible Medicaid expansion policies
(a) policy selected by the sample-analog rule
(b) policy selected by the trade-off rule
Notes: This figure plots the more flexible Medicaid expansion policies selected by statistical rules based on results from the OHIE. The horizontal dashed line marks the income thresholds under the current expansion policy, which is 138% regardless of the number of children in a household. The horizontal solid lines mark the more flexible policy selected by various statistical rules, i.e. income thresholds that can vary with number of children. For each number of children, I also plot the underlying income distribution to visualize individuals below the thresholds. Panel (a) plots the policy selected by the sample-analog rule $\widehat{g}_{\text{sample}}$. Panel (b) plots the policy selected by the trade-off rule $\widehat{g}_{\text{tradeoff}}$.
@@ Appendix A Proofs of theorems
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\left\{\omega:B(\widehat{g}(\omega);P)>k\right\}$ $>0.$ (19)
@@
$\limsup_{n\rightarrow\infty}Pr_{P_{0}^{n}}\left\{\omega:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega);P_{0})>\epsilon\right\}=0.$ (20)
Asymptotical feasibility under $P_{0}$ implies $Pr_{P_{0}^{n}}\{\omega:B(\widehat{g}(\omega);P_{0})\leq k\}\rightarrow 1.$
Consider the event $\omega^{\prime}$ where $\left|W(\widehat{g}(\omega^{\prime});P)-W(g_{P}^{\ast};P)\right|<\epsilon$ and $\widehat{g}(\omega^{\prime})$ is feasible. Asymptotic welfare-efficiency and asymptotic feasibility imply $Pr_{P_{0}^{n}}\left\{\omega^{\prime}\right\}\rightarrow 1.$ To see this, note the probability of such event has an asymptotic lower bound of one
$Pr_{P_{0}^{n}}\left\{\omega^{\prime}:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega^{\prime});P_{0})\leq\epsilon\ and\ B(\widehat{g}(\omega^{\prime});P_{0})\leq k\right\}$
$\geq$ $Pr_{P_{0}^{n}}\left\{\omega:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega);P_{0})\leq\epsilon\right\}+Pr_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})\leq k\right\}-1$
@@
$Pr_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})=k\right\}\geq Pr_{P_{0}^{n}}\left\{\omega^{\prime}\right\}\cdot Pr_{P_{0}^{n}}\left\{B(\widehat{g};P_{0})=k\mid\omega^{\prime}\right\}$
Then the above argument shows $Pr_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})=k\right\}\rightarrow 1$ as $n\rightarrow\infty$.
@@
$\mathcal{G}_{0}=\{g\in\mathcal{G}:B(g;P_{0})=k\}.$ (21)
Under Assumption 2.1, the sequence $P_{h_{n}}^{n}$ is contiguous with respect to the sequence $P_{0}^{n}$ , which means $P_{0}^{n}(A_{n})\rightarrow 0$ implies $P_{h_{n}}^{n}(A_{n})\rightarrow 0$ for every sequence of measurable sets $A_{n}$ on $\mathcal{A}^{n}$. Then $Pr_{P_{0}^{n}}\left\{A_{n}:B(\widehat{g}(A_{n});P_{0})=k\right\}\rightarrow 1$ implies there exists an $N(u)$ such that for all $n\geq N(u)$, we have $Pr_{P_{h_{n}}^{n}}\left\{A_{n}:B(\widehat{g}(A_{n});P_{0})=k\right\}\geq 1-u.$ That is, with high probability, the statistical rule $\widehat{g}$ selects policies from $\mathcal{G}_{0}$ based on the observed sample distributed according to $P_{h_{n}}^{n}$. Recall Assumption 2.1 implies for all $g\in\mathcal{G}_{0}$, for any sample size $n$, we have $B(g;P_{h_{n}})-k>C/\sqrt{n}.$ Thus this statistical rule cannot uniformly satisfy the constraint since with sample size $n\geq N(u)$, we have
$\sup_{P\in\mathcal{P}}Pr_{P^{n}}\left\{B(\widehat{g};P)>k\right\}\geq Pr_{P_{h_{n}}^{n}}\left\{B(\widehat{g};P_{h_{n}})>k\right\}\geq 1-u$ (22)
@@
Proof of Proposition 3.1. Consider a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$, which includes thresholds for a one-dimensional continuous characteristic $X$.
Suppose the budget constraint takes the form of a capacity constraint, involving a binary take-up decision $R$ (that may or may not be independent of $X$). By assumption, the probability an individual takes up the treatment for $X\in[\underline{t},\overline{t}]$ is zero but between zero and one otherwise. Then the budget function $B(t;P)\coloneqq E_{P}[R\cdot\mathbf{1}\{X\leq t\}]$ is flat in the interval $[\underline{t},\overline{t}]$ but also strictly increasing otherwise. Since $\Gamma>0$ almost surely, the welfare function $W(t;P)\coloneqq E_{P}[\Gamma\cdot\mathbf{1}\{X\leq t\}]$ is strictly increasing in $t$ and the constrained optimal policy satisfies Assumption 2.2 and is binding. Let $\inf_{B(g;P)<k}W(g_{P}^{\ast};P)-W(g;P)=E_{P}[\Gamma\cdot\mathbf{1}\{\underline{t}\leq X\leq\overline{t}\}]=\epsilon$ be the smallest amount of welfare loss from missing the binding solution $g_{P}^{\ast}$. Then $\epsilon>0$ and satisfies Assumption 2.3.
@@
$\max_{t}\ W(t;P)\ s.t.\ B(t;P)\leq k,$
where $B(t;P)=k$ for $t\in[\underline{t},\overline{t}]$ by assumption. The constrained optimal threshold is therefore the highest threshold where the constraint is satisfied exactly i.e. $t^{\ast}=\overline{t}.$ This also implies $R\cdot\mathbf{1}\{X\leq t\}\sim Bernoulli(k)$ for $t\in[\underline{t},\overline{t}]$.
@@
$\max_{t}\frac{1}{n}\sum_{i}\Gamma_{i}\cdot\mathbf{1}\{X_{i}\leq t\}\ s.t.\ \widehat{B}_{n}(t)\leq k$
for $\widehat{B}_{n}(t)\coloneqq\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq t\}$. Given that $\Gamma>0$ almost surely, the sample-analog rule equivalently solves $\max_{t}\widehat{B}_{n}(t)\ s.t.\ \widehat{B}_{n}(t)\leq k$. However, the solution is not unique because $\widehat{B}_{n}(t)$ is a step function. To be conservative, let the sample-analog rule be the smallest possible threshold to maximize $\widehat{B}_{n}(t)$:
$\widehat{t}=\min\left\{\arg\max_{t}\{\widehat{B}_{n}(t)\ s.t.\ \widehat{B}_{n}(t)\leq k\}\right\}.$
Note that we can also write $\widehat{B}_{n}(t)=\frac{1}{n}\sum_{R_{i}=1}\mathbf{1}\{X_{i}\leq t\}$, which makes it clear that $\widehat{t}$ corresponds to ranking $X_{i}$ among individuals with $R_{i}=1$, and then picking the lowest threshold such that we assign treatment to the first $\left\lfloor k\cdot n\right\rfloor$ individuals. This also means if in the sample few individuals take up the treatment such that $\frac{1}{n}\sum_{i}R_{i}\leq k$, we can have a sample-analog rule that treats everyone up to $\max_{R_{i}=1}X_{i}$. Taken together both scenarios, we note the sample-analog rule implies the treated share in the sample is equal to
$\widehat{B}_{n}(\widehat{t})\coloneqq\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq\widehat{t}\}=\min\left\{\frac{1}{n}\sum_{i}R_{i},\frac{\left\lfloor k\cdot n\right\rfloor}{n}\right\}.$
@@
$Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\widehat{B}_{n}(\widehat{t})\right\}$
$=$ $Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\ and\ \frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}R_{i}\right\}$
$+Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{1}{n}\sum_{i}R_{i}\ and\ \frac{\left\lfloor k\cdot n\right\rfloor}{n}\geq\frac{1}{n}\sum_{i}R_{i}\right\}$ (23)
$\geq$ $Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\ and\ \frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}R_{i}\right\}$
$\geq$ $Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\right\}+Pr_{P^{n}}\left\{\frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}R_{i}\right\}-1$ (24)

### v3

@@ 5.1 Simulation results
  • trade-off rule $\widehat{g}_{\text{tradeoff}}$ with $\bar{\lambda}=1$ and $r=\mathdollar 60,000$, same as the empirical illustration as explained in Equation (13).
@@
$\widehat{g}_{\text{sample}}$ $\widehat{g}_{\alpha=5\%}$ $\widehat{g}_{\text{tradeoff}}$
Prob. of selecting infeasible policies 10.2% 0.2% 45.2%
Prob. of selecting suboptimal policies 94.0% 99.4% 65.6%
Average welfare loss 0.06 0.17 0.01
Average budget overrun -\$70 -\$278 \$62
Notes: This table reports properties of statistical rules $\widehat{g}$, as averaged over 500 simulations. Row 1 reports the probability that the rule selects an eligibility policy that violates the budget constraint, i.e. $\operatorname{Pr}_{P^{n}}\{B(\widehat{g};P)>0\}$. Row 2 reports the probability that the rule achieves strictly less welfare than the constrained optimal policy $g_{P}^{\ast}$, i.e. $\operatorname{Pr}_{P^{n}}\{W(\widehat{g};P)<W(g_{P}^{\ast};P)\}$. Row 3 reports the average welfare loss of the rule relative to the maximum feasible welfare, i.e. $\frac{E_{P^{n}}\left[W(g_{P}^{\ast};P)-W(\widehat{g};P)\right]}{W(g_{P}^{\ast};P)}$. Row 4 reports the average budget overrun of the policies selected by the rule, i.e. $E_{P^{n}}\left[B(\widehat{g};P)\right]-k$.
Row 1 of Table 5.1 illustrates that it is possible for all three statistical rule $\widehat{g}$ to select infeasible policies. A lower probability of selecting infeasible policies suggests the rule is closer to achieving asymptotic feasibility. In the distribution calibrated to the OHIE sample, the original sample-analog rule $\widehat{g}_{\text{sample}}$ might not be asymptotically feasible as it can select infeasible eligibility policies in 10.2% of the draws. In contrast, Theorem 3.1 guarantees that a simple modification $\widehat{g}_{\alpha=5\%}$ selects infeasible eligibility policies in less than 5% of the draws, regardless of the distribution. Simulation confirms such guarantee as the mistakes only happen 0.2% of the time.
@@
In the distribution calibrated to the OHIE, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ on average achieves higher welfare than the sample-analog rule $\widehat{g}_{\text{sample}}$. As shown in row 3 of Table 5.1, the welfare loss of $\widehat{g}_{\text{tradeoff}}$ is 1% of the maximum feasible welfare $W(g_{P}^{\ast};P)$, compared to 6% for $\widehat{g}_{\text{sample}}$. However, its improvement can be at the cost of violating the budget constraint more often than $\widehat{g}_{\text{sample}}$, at a rate of 45.2%. Though as shown in row 4 of Table 5.1, on average the violation is limited, which confirms Theorem 4.1.
@@ Appendix A Proofs of theorems
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\left\{\omega:B(\widehat{g}(\omega);P)>k\right\}$ $>0.$
@@
$\limsup_{n\rightarrow\infty}\operatorname{Pr}_{P_{0}^{n}}\left\{\omega:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega);P_{0})>\epsilon\right\}=0.$
Asymptotical feasibility under $P_{0}$ implies $\operatorname{Pr}_{P_{0}^{n}}\{\omega:B(\widehat{g}(\omega);P_{0})\leq k\}\rightarrow 1.$
Consider the event $\omega^{\prime}$ where $\left|W(\widehat{g}(\omega^{\prime});P)-W(g_{P}^{\ast};P)\right|<\epsilon$ and $\widehat{g}(\omega^{\prime})$ is feasible. Asymptotic welfare-efficiency and asymptotic feasibility imply $\operatorname{Pr}_{P_{0}^{n}}\left\{\omega^{\prime}\right\}\rightarrow 1.$ To see this, note the probability of such event has an asymptotic lower bound of one
$\operatorname{Pr}_{P_{0}^{n}}\left\{\omega^{\prime}:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega^{\prime});P_{0})\leq\epsilon\text{ and }B(\widehat{g}(\omega^{\prime});P_{0})\leq k\right\}$
$\geq$ $\operatorname{Pr}_{P_{0}^{n}}\left\{\omega:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega);P_{0})\leq\epsilon\right\}+\operatorname{Pr}_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})\leq k\right\}-1$
@@
$\operatorname{Pr}_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})=k\right\}\geq\operatorname{Pr}_{P_{0}^{n}}\left\{\omega^{\prime}\right\}\cdot\operatorname{Pr}_{P_{0}^{n}}\left\{B(\widehat{g};P_{0})=k\mid\omega^{\prime}\right\}$
Then the above argument shows $\operatorname{Pr}_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})=k\right\}\rightarrow 1$ as $n\rightarrow\infty$.
@@
$\mathcal{G}_{0}=\{g\in\mathcal{G}:B(g;P_{0})=k\}.$
Under Assumption 2.1, the sequence $P_{h_{n}}^{n}$ is contiguous with respect to the sequence $P_{0}^{n}$ , which means $P_{0}^{n}(A_{n})\rightarrow 0$ implies $P_{h_{n}}^{n}(A_{n})\rightarrow 0$ for every sequence of measurable sets $A_{n}$ on $\mathcal{A}^{n}$. Then $\operatorname{Pr}_{P_{0}^{n}}\left\{A_{n}:B(\widehat{g}(A_{n});P_{0})=k\right\}\rightarrow 1$ implies there exists an $N(u)$ such that for all $n\geq N(u)$, we have $\operatorname{Pr}_{P_{h_{n}}^{n}}\left\{A_{n}:B(\widehat{g}(A_{n});P_{0})=k\right\}\geq 1-u.$ That is, with high probability, the statistical rule $\widehat{g}$ selects policies from $\mathcal{G}_{0}$ based on the observed sample distributed according to $P_{h_{n}}^{n}$. Recall Assumption 2.1 implies for all $g\in\mathcal{G}_{0}$, for any sample size $n$, we have $B(g;P_{h_{n}})-k>c/\sqrt{n}.$ Thus this statistical rule cannot uniformly satisfy the constraint since with sample size $n\geq N(u)$, we have
$\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\left\{B(\widehat{g};P)>k\right\}\geq\operatorname{Pr}_{P_{h_{n}}^{n}}\left\{B(\widehat{g};P_{h_{n}})>k\right\}\geq 1-u$
@@
Proof of Proposition 3.1.
By assumption, the probability an individual takes up the treatment for $X\in[\underline{t},\overline{t}]$ is zero but between zero and one otherwise. Then the budget function $B(t;P)\coloneqq\mathbb{E}_{P}[C\cdot\mathbf{1}\{X\leq t\}]$ is flat in the interval $\underline{t},\overline{t}]$ but strictly increasing otherwise. Since $\tau>0$ almost surely, the welfare function $W(t;P)\coloneqq\mathbb{E}_{P}[\tau\cdot\mathbf{1}\{X\leq t\}]$ is strictly increasing in $t$ and the constrained optimal policy satisfies Assumption 2.2 and is binding. Let $\inf_{B(g;P)<k}W(g_{P}^{\ast};P)-W(g;P)=\mathbb{E}_{P}[\tau\cdot\mathbf{1}\{\underline{t}\leq X\leq\overline{t}\}]=\epsilon$ be the smallest amount of welfare loss from missing the binding solution $g_{P}^{\ast}$. Then $\epsilon>0$ and satisfies Assumption 2.3.
@@
$\max_{t}\ W(t;P)\text{ subject to }B(t;P)\leq k,$
where $B(t;P)=k$ for $t\in[\underline{t},\overline{t}]$ by assumption. The constrained optimal threshold is therefore the highest threshold where the constraint is satisfied exactly i.e. $t^{\ast}=\overline{t}.$ This also implies $C\cdot\mathbf{1}\{X\leq t\}\sim Bernoulli(k)$ for $t\in[\underline{t},\overline{t}]$.
@@
$\max_{t}\frac{1}{n}\sum_{i}\tau_{i}\cdot\mathbf{1}\{X_{i}\leq t\}\text{ subject to }\widehat{B}_{n}(t)\leq k$
for $\widehat{B}_{n}(t)\coloneqq\frac{1}{n}\sum_{i}C_{i}\cdot\mathbf{1}\{X_{i}\leq t\}$. Given that $\tau>0$ almost surely, the sample-analog rule equivalently solves $\max_{t}\widehat{B}_{n}(t)\text{ subject to }\widehat{B}_{n}(t)\leq k$. However, the solution is not unique because $\widehat{B}_{n}(t)$ is a step function. To be conservative, let the sample-analog rule be the smallest possible threshold to maximize $\widehat{B}_{n}(t)$:
$\widehat{t}=\min\left\{\arg\max_{t}\{\widehat{B}_{n}(t)\text{ subject to }\widehat{B}_{n}(t)\leq k\}\right\}.$
Note that we can also write $\widehat{B}_{n}(t)=\frac{1}{n}\sum_{C_{i}=1}\mathbf{1}\{X_{i}\leq t\}$, which makes it clear that $\widehat{t}$ corresponds to ranking $X_{i}$ among individuals with $C_{i}=1$, and then picking the lowest threshold such that we assign treatment to the first $\left\lfloor k\cdot n\right\rfloor$ individuals. This also means if in the sample few individuals take up the treatment such that $\frac{1}{n}\sum_{i}C_{i}\leq k$, we can have a sample-analog rule that treats everyone up to $\max_{C_{i}=1}X_{i}$. Taken together both scenarios, we note the sample-analog rule implies the treated share in the sample is equal to
$\widehat{B}_{n}(\widehat{t})\coloneqq\frac{1}{n}\sum_{i}C_{i}\cdot\mathbf{1}\{X_{i}\leq\widehat{t}\}=\min\left\{\frac{1}{n}\sum_{i}C_{i},\frac{\left\lfloor k\cdot n\right\rfloor}{n}\right\}.$
@@
[…]
$=$ $\operatorname{Pr}_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\text{ and }\frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}C_{i}\right\}$
$+\operatorname{Pr}_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{1}{n}\sum_{i}C_{i}\text{ and }\frac{\left\lfloor k\cdot n\right\rfloor}{n}\geq\frac{1}{n}\sum_{i}C_{i}\right\}$
$\geq$ $\operatorname{Pr}_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\text{ and }\frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}C_{i}\right\}$
$\geq$ $\operatorname{Pr}_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\right\}+\operatorname{Pr}_{P^{n}}\left\{\frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}C_{i}\right\}-1$ (14)